Write a balanced chemical equation for the production of CaO(s) from the decomposition reaction of

limestone, CaCO3(s), into carbon dioxide gas and solid calcium oxide. Make sure to include the states

of each substance in the reaction.


Neutralization of Lake Weslemkoon near Bancroft, Ontario requires 10.0 kg of CaO. If you have 12.0

kg of CaCO3 available in your lab, do you have enough CaCO3 to produce the amount of CaO

necessary to neutralize the lake?


1
Expert's answer
2021-04-14T06:11:31-0400

CaCO3(s) = CaO(S) + CO2(g)


CaCO3(s) = CaO(S)

100g CaCO3(s) = 56g CaO(S)

1g CaCO3(s) = 56/100g CaO(S)

12000g CaCO3(s) = 56 × 12000g/100 CaO(S)

= 6720g CaO(S)

= 6.72Kg which will not be enough